# Observational constraints on generalised axion-like potentials for the late Universe

### v1

@@ 1 Introduction
Besides, axion-inspired scalar field models have been widely explored in cosmology as candidates for both dark energy and dark matter, with applications spanning from early universe inflation to present-day cosmic acceleration. These models often involve potentials motivated by axion physics, which naturally arise in several fundamental theories. Among them, a widely studied class of models adopts potentials of the form $V(\phi)\propto\left[1-\cos{(\phi/\eta)}\right]^{-n}$. Such potentials emerge in high-energy frameworks and have been investigated in various cosmological scenarios. The case where $n<0$ has been extensively explored in the literature, including applications in the string axiverse as a solution to the coincidence problem [20, 21], in early dark energy models as a means to address the Hubble tension [22, 23], in Natural Inflation models [24], and in primordial black hole formation mechanisms [25]. Additionally, the same potential has been used in axionic dark matter studies [26, 27], where oscillations around the potential minimum play a crucial role in defining the dynamics of ultralight axions. Cosmological simulations involving such ultralight axions have been performed in [28, 29, 30], demonstrating their viability as dark matter candidates. For further discussions on axions in cosmology, see [31, 32, 33, 34].
In this work, we focus on a specific dark energy model within this broad axion-inspired framework: the generalised axion-like dark energy model, first introduced in [35] and further analysed in [36, 37, 38]. In contrast to the more commonly studied $n<0$ case, here we analyse the scenario where $n>0$. The terminology generalised axion-like stems from its resemblance to the axion potentials appearing in particle physics, with the key distinction that we consider a positive $n$. The studies presented in [35, 36] demonstrated that this potential exhibits tracking behaviour, making it a compelling alternative to other quintessence models. Tracking solutions allow the scalar field to dynamically adjust to the evolution of the dominant energy component in the universe, alleviating the coincidence problem. Additionally, this model provides a natural mechanism for a smoother transition from matter domination to dark energy, making it preferable over alternative quintessence potentials. Further comparisons with the inverse power-law potential were made in [35], where it was shown that the generalised axion-like potential leads to a more rapid transition from matter to dark energy while ensuring that the present-day equation of state remains closer to $w=-1$. These properties suggest that this model provides a promising alternative to both $\Lambda$CDM and traditional quintessence models.
@@ 2.1 Model description
where $k^{2}=8\pi G=1/M_{P}^{2}$ (we use natural units $\hbar=c=1$ throughout this work and $M_{P}\approx 2.435\times 10^{18}\,\textrm{GeV}$ is the reduced Planck mass). $R$ denotes the Ricci scalar, and $L_{m}$, $L_{r}$, and $L_{\phi}$ correspond to the Lagrangians for matter, radiation, and the scalar field, respectively. For the scalar field, we define the Lagrangian as
@@
where $\Lambda$, $\eta$, and $n$ are positive free parameters of the model.
@@
A key feature of our model is its tracking behaviour. To analyse this behaviour, we introduce the parameter $\lambda$ [66, 67, 68, 69]:
$\lambda=-\frac{V_{,\phi}}{kV},$ (2.8)
which for our potential becomes [36]
When $\lambda$ is initially large ($\lambda_{i}\gg 1$), the scalar field exhibits tracking behaviour [36]. We further characterise the potential’s curvature using the parameter $\Gamma$ [66, 68, 69]:
$\Gamma=\frac{VV_{,\phi\phi}}{V_{,\phi}^{2}},$ (2.10)
which for our potential becomes [36]
$\Gamma=1+\frac{1}{2n}+\frac{n}{2k^{2}\eta^{2}\lambda^{2}}.$ (2.11)
The tracking phase is sustained as long as $\Gamma>1$ and $\Gamma\approx const.$ [66]. Under these conditions, the scalar field equation of state remains slightly below that of the dominant background component. For example, during the matter-dominated epoch ($w=0$), we derive [36]
$w_{\phi}\approx-\frac{1}{1+n}.$ (2.12)
This behaviour is crucial for addressing the coincidence problem, as it ensures that the scalar field energy density remains relevant throughout the universe’s evolution, avoiding an unnatural fine-tuning of initial conditions.
$V(\phi)\approx V_{\rm min}+\frac{1}{2}m^{2}(\phi-\pi\eta)^{2},$ (2.13)
where we have defined $V_{\rm min}\equiv\Lambda^{4}/2^{n}$ and $m^{2}\equiv n\Lambda^{4}/(2^{n+1}\eta^{2})$. Note that we can use $V_{\rm min}$ and $\Lambda$ interchangeably as a free parameter of the model. In the following section, we will use $V_{\rm min}$ in the study of the observational constraints. At the end of this stage, the kinetic energy term becomes negligible, and the scalar field behaves like dark energy, fuelling accelerated cosmic expansion. This late-time attractor behaviour ensures the equation of state approaches $-1$. Unlike tracking, where $w_{\phi}$ depends on the parameter $n$, the field dominance at late-time results in a behaviour that closely mimics a cosmological constant. The rapid growth of $\Gamma$ in the final stages of the field evolution further accelerates this transition, distinguishing the generalised axion-like model from simpler inverse power-law potentials [35, 36].
2.2 Cosmological evolution
In this subsection, we present some results obtained from the numerical study of the dynamical system of the generalised axion-like model. A detailed and comprehensive analysis of this study was conducted in [36]. We next highlight the most important results for our current work.
The background evolution of a general quintessence model can be expressed through a closed autonomous system of differential equations given by
@@ 2.2 Cosmological evolution
where a prime denotes a derivative with respect to $\ln{(a/a_{0})}$ ), and $a_{0}$ is the current value of the scale factor. Here we have introduced the compactified variables $x$ and $y$, which correspond, respectively, to the kinetic and potential terms of the scalar field [67]:
@@
$z=\Omega_{r}^{1/2}=\frac{k\sqrt{\rho_{r}}}{\sqrt{3}H}.$ (2.19)
@@
$f(\lambda)=\lambda^{2}[\Gamma(\lambda)-1]=\frac{n}{2k^{2}\eta^{2}}+\frac{\lambda^{2}}{2n}.$ (2.20)
@@
In Fig. 1(a), we present the evolution of the scalar field equation of state, $w_{\phi}$, for the case $n=0.5$, $\eta=M_{P}$. The results confirm the tracking behaviour: following the frozen phase, where $w_{\phi}\approx-1$, all trajectories converge onto a common tracking path, regardless of their initial conditions. This universality in the evolution of $w_{\phi}$ demonstrates the attractor nature of the tracking solution. The equation of state in the tracking regime can be analytically determined using (2.12). For the specific case examined, $n=0.5$, we find $w_{\phi}\approx-2/3$. These theoretical predictions align well with the numerical results shown in the figure. A detailed analysis of the tracking behaviour in this model was carried out in [36].
@@ 3 Observational fit
Regarding the prior, we follow the standard practice of [7] for the early-time cosmological parameters, as shown in table 1. Additional BBN (big bang nucleosynthesis) prior on the baryon density $\Omega_{b0}h^{2}$ and a cut $10040<H_{0}<$ 100 on the Hubble constant are imposed across all models and datasets considered in this work to stabilise chains without CMB data. For late-time physics, i.e., the dark energy model, we put the axion-like dark energy model in section 2 with $\dot{\phi}(a_{i}/a_{0}=10^{-6})=0$ against the $\Lambda$CDM model. The constraint on the initial field momentum is reasonable given the freezing property during radiation dominated era demonstrated in section 2.
@@ 3.1 Data
We consider multiple astrophysical and cosmological datasets. For simplicity, they are grouped together according to the theoretical origin as CMB (Planck2018 $-\ell$ TTEE [39], NPIPE CamSpec $-\ell$ TTTEEE [16], PR4 CMB lensing [40, 41]) , BAO (DESI DR1 without full-shape [42, 43, 44]), SNe (Pantheon+ SNe catalog [18] without anchoring of SNe standardised absolute magnitude), low-$z$ (Riess et al. low-$z$ anchors [17]), and DES Y1 morphological datasets [45]. Additional $f\sigma_{8}$ dataset (table 2 of [76], with data entries from [51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63]), SDSS DR7 LRG dataset [64] and eBOSS DR14 Ly-$\alpha$ dataset [65] are presented in our figures in section 3.3, but are excluded from our cosmological fit.
@@ 3.2 Cosmological parameters
For the model parameters of the axion-like dark energy, we choose the excess dark energy density over the potential minimum $\log_{10}(\rho_{\rm DE,0}/V_{\rm min}-1)$ and the curvature of the potential around the minimum $\log_{10}(m^{2}/V_{\rm min})=\log_{10}(n\eta^{-2}/2)$ defined in Eq. (2.13) instead of $\log_{10}(V_{\rm min}/\rho_{\rm DE,0})$ and $\log_{10}(\eta/M_{P})$, as these two are directly tied to the late-time observables. The prior of $\log_{10}(\rho_{\rm DE,0}/V_{\rm min}-1)$ would be non-flat though, with a density that penalises negative values linearly in terms of log prior probability.
@@
Figure 3: 68% and 95% C.L. posterior distribution of the early-time parameters from CMB (Black), CMB + BAO (blue), CMB + BAO + SNe (red), CMB + BAO + SNe + low-$z$ ( green), and CMB + BAO + SNe + low-$z$ + DES Y1 (Brown) data, with $\Lambda$CDM model the lighter shade contours and the axion-like dark energy model in section 2 the darker shade contours. The result demonstrates the stability of the early-time physics while strongly indicates the inability of the axion-like dark energy model to ease the tension between astrophysical data of low-$z$ and cosmological data of CMB + BAO.
Figure 4: 68% and 95% C.L. posterior distribution of the late-time quantities from CMB (Black), CMB + BAO (blue), CMB + BAO + SNe (red), CMB + BAO + SNe + low-$z$ ( green), and CMB + BAO + SNe + low-$z$ + DES Y1 (Brown) data, with $\Lambda$CDM model the lighter shade contours and the axion-like dark energy model in section 2 the darker shade contours. The result confirms the inability of the axion-like dark energy model to increase the predicted $H_{0}$ value.
However, Figs. 4 only show the inability of the axion-like dark energy model to modify CMB predictions, which is half of the story. To confirm the persistence of the tension in the axion-like dark energy model, we perform the MCMC analyses of BAO, SNe+low-$z$ and DES Y1 datasets alone without CMB dataset. Mean and standard deviation of the cosmological parameters, late-time observables, and statistical probes are shown in table 4 for the $\Lambda$CDM model and in table 5 for the axion-like dark energy model. Notice that the lack of constraint on $\sigma_{8}$ and $S_{8}$ for BAO and SNe + low-$z$ datasets is expected, as we impose no a priori constraint on the primordial perturbation. Similarly, $n_{s}$ and $\tau_{\rm reio}$ are omitted due to the lack of constraint by the late-time physics. $\Omega_{b0}h^{2}$ is omitted as well since it would only reflect the BBN prior. In addition, the DES Y1 posterior for the axion-like dark energy model is a rough estimate, as we cannot reach the stopping criteria.
@@
Following the statistical analysis described at the beginning of section 3, we carry out the statistical probes of Bayesian ratio, goodness of fit and suspiciousness, and find that the axion-like dark energy model overall shows no statistical (dis)-advantage over $\Lambda$CDM model in terms of data tension.
However, as shown in Fig. 5, the $2$-dimensional posterior distribution shifts significantly from $\Lambda$CDM model to the axion-like dark energy model, especially for SNe + low-$z$ dataset. This may contribute to the very slight ease of tension between CMB + BAO dataset and SNe + low-$z$ dataset, as shown in table 5. It mainly comes from Eq. (2.12), where the increase of dark energy equation of state during the tracking regime leads to less dark matter at the current time than what would be expected from the luminosity distance measurements. Unfortunately, the Hubble tension remains despite the introduction of the dynamical field.
Figure 5: 68% and 95% C.L. posterior distribution of the late-time quantities from CMB (Black), BAO (blue), SNe + low-$z$ (red), and DES Y1 (Brown) data, with $\Lambda$CDM model the lighter shade contours and the axion-like dark energy model in section 2 the darker shade contours. The strong tension between astrophysical data of SNe + low-$z$ and cosmological data of CMB + BAO presented in $\Lambda$CDM model posterior of $H_{0}$ persists in the axion-like dark energy model. The lack of constraint on $S_{8}$ for BAO and SNe + low-$z$ datasets is expected as we impose no a priori constraint on the primordial perturbation. Note that the DES Y1 posterior is a rough estimate, as explained in section 3.1.
To better understand how the axion-like dark energy model affects the cosmological fit and why it cannot solve the Hubble tension, the extended parameters of the axion-like dark energy model are shown in Fig. 6 for SNe + low-$z$ dataset and Fig. 7 for CMB dataset as scatter plots over the current matter density parameter $\Omega_{m0}$. We choose to omit $H_{0}$ as it is either fully constrained (SNe + low-$z$) or is strongly correlated with $\Omega_{m0}$ (CMB) and thus provides no additional information.
@@
For Fig. 6, the effect of the axion-like dark energy on $\Omega_{m0}$ prediction of SNe + low-$z$ dataset is tremendous. This is the direct consequence of Eq. (2.12) that a larger scalar equation of state $w_{\phi}$ leads to a lower value of $\Omega_{m0}$ for a constant $H_{0}$. Furthermore, one can clearly identify the tracking regime in $\log(n\eta^{-2})$ $\log(\phi_{i}/\eta)$ scatter plot. Heavier fields ($\log(2m^{2}/V_{\rm min})=\log(n\eta^{-2})\gg 0$) roll down the potential faster, thus ending the tracking regime at higher redshift, affecting SNe + low-$z$ dataset prediction less. Lighter fields ($\log(n\eta^{-2})\ll 0$) on the other hand seems to be clustered around $n\sim 0$, suggesting that they are effectively $w$CDM with $w\sim-1$. To have non-trivial effect on SNe + low-$z$ prediction, the scalar field should exit the tracking regime during dark energy dominant era ($\log(n\eta^{-2})\sim 0$ and $\log(\phi_{i}/\eta)<0$) and should have distinguishable feature during the tracking regime ($n\nsim 0$).
For Fig. 7 though, the effect of the axion-like dark energy on $\Omega_{m0}$ prediction of CMB dataset is less pronounced. $\Omega_{m0}$ is pulled toward the opposite direction than the prediction of SNe + low-$z$ dataset. This is, again, consistent with Eq. (2.12) that a larger scalar equation of state $w_{\phi}$ (greater $H(z)/H_{0}$) necessitates a lower value of $H_{0}$ (effectively lower $\Omega_{m0}$ for a constant $\Omega_{m0}h^{2}$) to maintain a constant distance to the last scattering surface. The deviation from $w=-1$ is more effective at lower redshift as it has greater effect on the distance measure. A lighter field ($\log(n\eta^{-2})<2$) that stays long enough inside the tracking regime ($\log(\phi_{i}/\eta)<0$) thus has greater effect on $\Omega_{m0}$. Unfortunately, Fig. 7 is polluted by the finite width Gaussian constraint of $\mathcal{E}_{H_{0}}$ in table 1. The finite width causes an oversampling of heavier fields that have reached the bottom (the Gaussian peak at $\log_{10}(\rho_{\rm DE,0}/V_{\rm min}-1)\sim-3$), but would not affect the conclusion that the axion-like dark energy in section 2 predicts a lower value of $H_{0}$ for CMB dataset, thus unable to ease the Hubble tension.

### v2

@@ 1 Introduction
The observational challenges and opportunities call for theoretical investigation. In particular, axion-inspired scalar field models have been widely explored in cosmology as candidates for both dark energy and dark matter, with applications spanning from early universe inflation to present-day cosmic acceleration. These models often involve potentials motivated by axion physics, which naturally arise in several fundamental theories. Among them, a widely studied class of models adopts potentials of the form $V(\phi)\propto\left[1-\cos{(\phi/\eta)}\right]^{-n}$. Such potentials emerge in high-energy frameworks and have been investigated in various cosmological scenarios. The case where $n<0$ has been extensively explored in the literature, including applications in the string axiverse as a solution to the coincidence problem [20, 21], in early dark energy models as a means to address the Hubble tension [22, 23], in Natural Inflation models [24], and in primordial black hole formation mechanisms [25]. Additionally, the same potential has been used in axionic dark matter studies [26, 27], where oscillations around the potential minimum play a crucial role in defining the dynamics of ultralight axions. Cosmological simulations involving such ultralight axions have been performed in [28, 29, 30], demonstrating their viability as dark matter candidates. For further discussions on axions in cosmology, see [31, 32, 33, 34].
In this work, we focus on a specific dark energy model within this broad axion-inspired framework: the generalised axion-like dark energy model, first introduced in [35] and further analysed in [36, 37, 38]. In contrast to the more commonly studied $n<0$ case, here we analyse the scenario where $n>0$. The terminology generalised axion-like stems from its resemblance to the axion potentials appearing in particle physics, with the key distinction that we consider a positive $n$. The studies presented in [35, 36] demonstrated that this potential exhibits tracking behaviour, making it a compelling alternative to other quintessence models. Tracking solutions allow the scalar field to dynamically adjust to the evolution of the dominant energy component in the universe, alleviating the coincidence problem. Additionally, this model provides a mechanism for a smoother transition from matter domination to dark energy, making it preferable over alternative quintessence potentials. Further comparisons with the inverse power-law potential were made in [35], where it was shown that the generalised axion-like potential leads to a more rapid transition from matter to dark energy while ensuring that the present-day equation of state remains closer to $w=-1$. These properties suggest that this model provides a promising alternative to both $\Lambda$CDM and traditional quintessence models, by substituting the cosmological constant with a dynamical field that is free from additional fine-tuning.
@@ 2.1 Model description
where $k^{2}=8\pi G=1/M_{P}^{2}$ (we use natural units $\hbar=c=1$ throughout this work and $M_{P}\approx 2.435\times 10^{18}\,\textrm{GeV}$ is the reduced Planck mass). $R$ denotes the Ricci scalar, and $L_{m}$, $L_{r}$, and $L_{\phi}$ correspond to the Lagrangian densities for matter, radiation, and the scalar field, respectively. For the scalar field, we define the Lagrangian density as
@@
where $\Lambda$, $\eta$, and $n$ are positive free parameters of the model. In standard axion-like scenarios, $\eta$ is typically associated with the spontaneous symmetry-breaking scale of a global $U(1)$ symmetry, while $\Lambda$ represents the scale of non-perturbative effects that generate the potential. The exponent $n$ parametrises deviations from the usual cosine potential.
@@
The background evolution of this model can be fully described using a dynamical system approach. By introducing suitable compact variables, the equations of motion can be recast into a closed autonomous system. This allows us to identify critical points, study their stability, and characterise attractor solutions (see table 1). In particular, the model dynamics can be described by the following set of differential equations [36]:
@@
where a prime denotes a derivative with respect to $\ln(a/a_{0})$, and $a_{0}$ is the current value of the scale factor. Note that this set of equations corresponds to a reformulation of the original Friedmann and Klein–Gordon equations in terms of dimensionless variables, and rewritten as a first-order autonomous system. The variables $x$ and $y$ correspond, respectively, to the dimensionless kinetic and potential energy contributions of the scalar field [66]:
@@
$z=\Omega_{r}^{1/2}=\frac{k\sqrt{\rho_{r}}}{\sqrt{3}H},$ (2.13)
and the parameter $\lambda$ [67, 66, 68, 69]:
$\lambda=-\frac{V_{,\phi}}{kV},$ (2.14)
which for our potential becomes [36]
@@
$f(\lambda)=\lambda^{2}[\Gamma(\lambda)-1]=\frac{n}{2k^{2}\eta^{2}}+\frac{\lambda^{2}}{2n},$ (2.16)
where the parameter $\Gamma$ represents the potential’s curvature and is defined as [67, 68, 69]
$\Gamma=\frac{VV_{,\phi\phi}}{V_{,\phi}^{2}},$ (2.17)
which for our potential becomes [36]
$\Gamma=1+\frac{1}{2n}+\frac{n}{2k^{2}\eta^{2}\lambda^{2}}.$ (2.18)
 Point $x$   $y$ $z$ $\lambda$ $w_{\phi}$ Stability
A 00 00 $1$ $\lambda_{\star}$  Undefined Saddle
B 00 00 00 $\lambda_{\star}$  Undefined Saddle
C 00 $1$ 00 00 $-1$  Stable if $n>0$
D $x_{\star}$ 00 $z_{\star}$  Undefined $1$ Saddle if $n>0$
Table 1: Fixed points of the system given by (2.8), (2.9), (2.10) and (2.11). D is deduced by compactifying the $\lambda$ variable. In fact, $\lambda$ blows-up in this case (for more details, see [36]). The quantities $x_{\star}$, $y_{\star}$ and $\lambda_{\star}$ are random values of the variables $x$, $y$ and $\lambda$ inside their domains which satisfy the equations of motion.
The system admits a set of fixed points whose properties provide insight into the qualitative dynamics of the model. These fixed points, summarised in table 1, correspond to distinct cosmological regimes: points A and B are associated with radiation and matter domination, respectively, and generally appear as saddle points in dynamical analyses of quintessence models. Point C is a potential-dominated fixed point corresponding to a de Sitter state and describes the late-time dark energy-dominated regime. Unlike A and B, the stability of point C depends on the form of the potential. In our model, it is stable for $n>0$, and hence acts as the true attractor governing the asymptotic cosmic acceleration. Point D, by contrast, represents a kinetically dominated phase and arises in the regime where $\lambda\rightarrow\infty$. To consistently incorporate it into the phase space analysis, the variable $\lambda$ must be compactified (see [36] for more details).
A key feature of our model is the tracking behaviour, which appears only if $n>0$. This phase is associated with initial values $\lambda_{i}\gg 1$ (where the subscript $i$ denotes the initial condition hereafter), where the field slowly rolls down a steep potential slope and the curvature parameter $\Gamma$ remains nearly constant. In this regime, the system evolves along a family of instantaneous critical points in the $(x,y,z,\lambda)$ phase space, solutions that are not true fixed points, but which act as quasi-attractors guiding the field evolution [69]. This behaviour can be sustained as long as $\Gamma>1$ and $\Gamma\approx\text{const.}$. [67]. Under these conditions, the scalar field equation of state remains below that of the dominant component. For instance, during matter domination ($w=0$), it takes the approximate value [36]
$w_{\phi}\approx-\frac{1}{1+n}.$ (2.19)
This tracking mechanism allows the scalar field to evolve in a manner largely independent of its initial conditions.
$V(\phi)\approx V_{\rm min}+\frac{1}{2}m^{2}(\phi-\pi\eta)^{2},$ (2.20)
where we have defined $V_{\rm min}\equiv\Lambda^{4}/2^{n}$ and $m^{2}\equiv n\Lambda^{4}/(2^{n+1}\eta^{2})$. Note that we can use $V_{\rm min}$ and $\Lambda$ interchangeably as a free parameter of the model. In the following section, we will use $V_{\rm min}$ in the study of the observational constraints. On the other hand, the scalar mass $m$ directly controls the ending of the tracking phase. For the tracking to last for $O(H_{0}^{-1})$, we need $m^{2}/H_{0}^{2}\sim nM_{P}^{2}\eta^{-2}\sim O(1)$. The field strength $\phi$ thus potentially reaches super-Planckian scale. However, it does not immediately lead to trans-Planckian problem, as $\phi$ by itself is not an observable from the gravity point of view. At the end of this stage, the kinetic energy term becomes negligible, and the scalar field behaves like dark energy, fuelling accelerated cosmic expansion. In terms of the dynamical system, this final regime corresponds to the approach toward a true critical point: the de Sitter attractor $(x,y,z,\lambda)=(0,1,0,0)$, listed as point C in table 1. This point is stable for $n>0$ and represents a late-time universe with $w_{\phi}=-1$. Unlike the tracking phase, where $w_{\phi}$ depends on $n$, the de Sitter attractor yields a behaviour indistinguishable from a cosmological constant. The rapid growth of $\Gamma$ as the field approaches the minimum ensures the system settles quickly into this final state, distinguishing the generalised axion-like model from simpler inverse power-law potentials [35, 36].
2.2 Cosmological evolution
In this subsection, we present some results obtained from the numerical study of the dynamical system of the generalised axion-like model given by (2.8), (2.9), (2.10) and (2.11). A detailed and comprehensive analysis of this study was conducted in [36]. We next highlight the most important results for our current work.
@@ 2.2 Cosmological evolution
In Fig. 1(a), we present the evolution of the scalar field equation of state, $w_{\phi}$, for the case $n=0.5$, $\eta=M_{P}$. The results confirm the tracking behaviour: following the frozen phase, where $w_{\phi}\approx-1$, all trajectories converge onto a common tracking path, regardless of their initial conditions. This universality in the evolution of $w_{\phi}$ demonstrates the attractor behaviour of the tracking solution. The equation of state in the tracking regime can be analytically determined using (2.19). For the specific case examined, $n=0.5$, we find $w_{\phi}\approx-2/3$. These theoretical predictions align well with the numerical results shown in the figure. A detailed analysis of the tracking behaviour in this model was carried out in [36].
@@ 3 Observational fit
Regarding the prior, we follow the standard practice of [7] for the early-time cosmological parameters, as shown in table 2. Additional BBN (big bang nucleosynthesis) prior on the baryon density $\Omega_{b0}h^{2}$ and a cut $10040<H_{0}<$ 100 on the Hubble constant are imposed across all models and datasets considered in this work to stabilise chains without CMB data. For late-time physics, i.e., the dark energy model, we put the axion-like dark energy model in section 2 with $\dot{\phi}(a_{i}/a_{0}=10^{-6})=0$ (abbreviated as $\phi$CDM in figures) against the $\Lambda$CDM model. The constraint on the initial field momentum is reasonable given the freezing property during radiation dominated era demonstrated in section 2.
@@ 3.1 Data
We consider multiple astrophysical and cosmological datasets. For simplicity, they are grouped together according to the theoretical origin as CMB (Planck2018 $-\ell$ TTEE [39], NPIPE CamSpec $-\ell$ TTTEEE [16], PR4 CMB lensing [40, 41]) , BAO (DESI DR1 without full-shape [42, 43, 44]), SNe (Pantheon+ SNe catalog [18] without anchoring of SNe standardised absolute magnitude), low-$z$ (Riess et al. low-$z$ anchors [17]), and DES Y1 morphological datasets [45]. Additional $f\sigma_{8}$ dataset (table 2 of [76], with data entries from [51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63]), SDSS DR7 LRG dataset [64] and eBOSS DR14 Ly-$\alpha$ dataset [65] are presented in Figs. 8(e), 8(f) and 11 in section 3.3 for visual comparison. They are not considered in the MCMC analysis to avoid potential overlap with the DESI BAO dataset.
@@ 3.2 Cosmological parameters
For the model parameters of the axion-like dark energy, instead of the field-theory-inspired flat prior space of $\log_{10}(V_{\rm min}/\rho_{\rm DE,0})$ and $\log_{10}(\eta/M_{P})$, we re-parametrise them as the excess dark energy density over the potential minimum $\log_{10}(\rho_{\rm DE,0}/V_{\rm min}-1)$ and the curvature of the potential around the minimum $\log_{10}(m^{2}/V_{\rm min})=\log_{10}(n\eta^{-2}/2)$ defined in Eq. (2.20), as these two are directly tied to the late-time observables. Effectively, we arrive at a prior probability density in $\log_{10}(\rho_{\rm DE,0}/V_{\rm min}-1)$ space that is exponentially suppressed toward the negative direction.
@@
Figure 3: 68% and 95% C.L. posterior distribution of the early-time parameters from CMB (Black), CMB + BAO (blue), CMB + BAO + SNe (red), CMB + BAO + SNe + low-$z$ ( green), and CMB + BAO + SNe + low-$z$ + DES Y1 (Brown) data, with $\Lambda$CDM model the dashed contours and the axion-like dark energy model ($\phi$CDM) the solid contours. The result demonstrates the stability of the early-time physics.
Figure 4: 68% and 95% C.L. posterior distribution of the late-time quantities from CMB (Black), CMB + BAO (blue), CMB + BAO + SNe (red), CMB + BAO + SNe + low-$z$ ( green), and CMB + BAO + SNe + low-$z$ + DES Y1 (Brown) data, with $\Lambda$CDM model the dashed contours and the axion-like dark energy model ($\phi$CDM) the solid contours. The result confirms the inability of the axion-like dark energy model to increase the predicted $H_{0}$ value. The non-Gaussian shape of the contours is a feature of the model, as explained in appendix B.
However, the main goal of the model is not to alter the CMB prediction significantly, but to provide a dynamical scalar alternative to the cosmological constant. In this regard, we perform the MCMC analyses of BAO, SNe+low-$z$ and DES Y1 datasets alone without CMB dataset. Mean and standard deviation of the cosmological parameters, late-time observables, and statistical probes are shown in table 5 for the $\Lambda$CDM model and in table 6 for the axion-like dark energy model. Notice that the lack of constraint on $\sigma_{8}$ and $S_{8}$ for BAO and SNe + low-$z$ datasets is expected, as we impose no a priori constraint on the primordial perturbation. Similarly, $n_{s}$ and $\tau_{\rm reio}$ are omitted due to the lack of constraint by the late-time physics. $\Omega_{b0}h^{2}$ is omitted as well since it would only reflect the BBN prior. In addition, the DES Y1 posterior for the axion-like dark energy model is a rough estimate, as we cannot reach the stopping criteria.
@@
Following the statistical analysis described at the beginning of section 3, we carry out the statistical probes of Bayesian ratio, goodness of fit and suspiciousness, and find that the axion-like dark energy model overall shows no statistical (dis)-advantage over $\Lambda$CDM model in terms of data tension, demonstrating the potential of the dynamical, axion-like dark energy model as an alternative to the cosmological constant.
Yet, as shown in Fig. 5, the $2$-dimensional posterior distribution shifts significantly from $\Lambda$CDM model to the axion-like dark energy model, especially for SNe + low-$z$ dataset. This may contribute to the very slight ease of tension between CMB + BAO dataset and SNe + low-$z$ dataset, as shown in table 6. This is a direct consequence of Eq. (2.19), where the increase of dark energy equation of state during the tracking regime leads to less dark matter at the current time than what would be expected from the luminosity distance measurements, and is common among dark energy models with an equation of state deviating from -1.
Figure 5: 68% and 95% C.L. posterior distribution of the late-time quantities from CMB (Black), BAO (blue), SNe + low-$z$ (red), and DES Y1 (Brown) data, with $\Lambda$CDM model the dashed contours and the axion-like dark energy model ($\phi$CDM) the solid contours. The lack of constraint on $S_{8}$ for BAO and SNe + low-$z$ datasets is expected as we impose no a priori constraint on the primordial perturbation. Note that the DES Y1 posterior is a rough estimate, as explained in section 3.1.
To better understand how the axion-like dark energy model affects the cosmological fit, the extended parameters of the axion-like dark energy model are shown in Fig. 6 for SNe + low-$z$ dataset and Fig. 7 for CMB dataset as scatter plots over the current matter density parameter $\Omega_{m0}$. We choose to omit $H_{0}$ as it is either fully constrained (SNe + low-$z$) or is strongly correlated with $\Omega_{m0}$ (CMB) and thus provides no additional information.
@@
For Fig. 6, the effect of the axion-like dark energy on $\Omega_{m0}$ prediction of SNe + low-$z$ dataset is tremendous. Again, this is the direct consequence of Eq. (2.19) that a larger scalar equation of state $w_{\phi}$ leads to a lower value of $\Omega_{m0}$ for a constant $H_{0}$. Furthermore, one can clearly identify the tracking regime in $\log(n\eta^{-2})$ $\log(\phi_{i}/\eta)$ scatter plot. Heavier fields ($\log(2m^{2}/V_{\rm min})=\log(n\eta^{-2})\gg 0$) roll down the potential faster, thus ending the tracking regime at higher redshift, affecting SNe + low-$z$ dataset prediction less. Lighter fields ($\log(n\eta^{-2})\ll 0$) on the other hand seems to be clustered around $n\sim 0$, suggesting that they are effectively $w$CDM with $w\sim-1$. To have non-trivial effect on SNe + low-$z$ prediction, the scalar field should exit the tracking regime during dark energy dominant era ($\log(n\eta^{-2})\sim 0$ and $\log(\phi_{i}/\eta)<0$) and should have distinguishable feature during the tracking regime ($n\nsim 0$).
For Fig. 7 though, the effect of the axion-like dark energy on $\Omega_{m0}$ prediction of CMB dataset is less pronounced. $\Omega_{m0}$ is pulled toward the opposite direction than the prediction of SNe + low-$z$ dataset. This is, again, consistent with Eq. (2.19) that a larger scalar equation of state $w_{\phi}$ (greater $H(z)/H_{0}$) necessitates a lower value of $H_{0}$ (effectively lower $\Omega_{m0}$ for a constant $\Omega_{m0}h^{2}$) to maintain a constant distance to the last scattering surface. The deviation from $w=-1$ is more effective at lower redshift as it has greater effect on the distance measure. A lighter field ($\log(n\eta^{-2})<2$) that stays long enough inside the tracking regime ($\log(\phi_{i}/\eta)<0$) thus has greater effect on $\Omega_{m0}$. Unfortunately, Fig. 7 is polluted by the finite width Gaussian constraint of $\mathcal{E}_{H_{0}}$ in table 2. The finite width causes an oversampling of heavier fields that have reached the bottom (the Gaussian peak at $\log_{10}(\rho_{\rm DE,0}/V_{\rm min}-1)\sim-3$), but would not affect the conclusion that the axion-like dark energy predicts a lower value of $H_{0}$ for CMB dataset than the $\Lambda$CDM prediction if the tracking regime overlaps with the dark-energy-dominated era. The inclusion of the low-$z$ dataset eliminates this possibility, forcing the axion-like dark energy model to follow closely with $\Lambda$CDM. The oversampling from the finite width of $\mathcal{E}_{H_{0}}$ thus merely alters the posterior of the extended parameters and would not affect the result. For a more detailed analysis, see appendix B.